OpenStudy (anonymous):

We have a soln of protein that has a concentration of 0.25mg/ml. We need 20 ug of the protein for an experiment. What volume of the protein soln do we need?

OpenStudy (anonymous):

I know I have to write a conversion, I'm just confused at where I put what... 20ug x (____uL/ 0.25 ug) = _____? AM I close at all?

OpenStudy (anonymous):

m(in mg) * 0.001 = m(in ug) (1mg = 1000ug) m(in mg) = C(mg/ml) * V(in ml) so, 20ug * 0.001mg/ug = 0.02 mg 0.02mg = 0.25mg/ml * V V = 0.02/0.25 = 0.08ml

OpenStudy (anonymous):

THANK YOU SO MUCH!!! I needed help so bad!! You are awesome!!